# Using Large Language Models to Assist Video Content Analysis: An Exploratory Study of Short Videos on Depression

### v2

@@ Abstract.
Multimodal Information, Large Language-and-Vision Assistant (LLaVA), User Generated Content, Content Analysis, Mental Health
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: CSCW 24; ; booktitle: CSCW 24**footnotetext: These authors contributed equally to this work
@@ 1   Introduction
To address this gap, we propose a LLM-assisted content analysis workflow (Figure 1) that can both save human labor in the analysis and capture the nuances in the content. Recently, LLMs have been tested in various types of qualitative data analysis (Shen et al., 2023), such as deductive thematic analysis (Zhang et al., 2023); research findings show that LLMs are able to improve coding efficiency (Chew et al., 2023), facilitate collaborative coding (Gao et al., 2023), and increase inter-coder reliability (Xiao et al., 2023). However, most of the research is based on text data. In this paper, we join the exploration of LLMs for data analysis and situate it in the video-data context. Using LLMs, we are specifically interested in customizing codebooks and designing prompts to comprehend and capture more fine-grained information to encompass the semantic and contextual information conveyed in the video content.
@@ 2   Context and Dataset: Depression-
We plan to use LLMs to analyze the content of YouTube videos related to users’ self-disclosure of depression. Online self-disclosure, defined as sharing sensitive personal information in an online public space (Misoch, 2015), has been extensively studied on text-based platforms such as Twitter (Lachmar et al., 2017), Weibo (Wang et al., 2019), and Reddit (Chen and Link to external site, 2021). The multimodal nature of video may enable deeper levels of disclosure and facilitate a stronger sense of connection between creators and their audience (Liu and Zhang, 2023). By applying LLMs to extract quantifiable information from the videos automatically, we aim to enable future studies of exploring how people convey information and emotion via self-disclosure videos (Chaudoir and Fisher, 2010). This case study only uses this flow as a preliminary method to analyze video content to explore the capability and limitations of LLMs in content analysis.
@@ 5.1. Future Opportunities and Challenges
LLMs could reduce the human labor required for manual coding in video content analysis. Researchers can create a structured framework (e.g., a codebook) to guide LLMs to automatically identify and extract meaningful insights from unstructured video data. In our case, operationalizing the concept of depth of self-disclosure into various observable behaviors in different settings, such as the presence of other people and the behaviors disclosed, allows us to understand the nuances of how individuals express themselves in video content. However, we find the most critical step in this process is operationalizing concepts into well-defined codes, as it sets the foundation for accurate and comprehensive analysis by LLMs. Further, it is difficult for researchers to curate prompts to harness the power of LLMs (e.g., annotating abstract concepts such as genre in images, as shown in our case study) and effectively evaluate the structured output generated by LLMs. Future work can investigate how to better involve humans in the workflow.

### v3

# Harnessing LLMs for Automated Video Content Analysis: An Exploratory Workflow of Short Videos on Depression

@@ Abstract.
Multimodal Information; Visual Content; Images; Large Language-and-Vision Assistant (LLaVA); User Generated Content; Large Language Model; Mental Health
$c$s: Computing methodologiesccs: Computing methodologies Computer visionccs: Human-centered computing Collaborative and social computing design and evaluation methodsccs: Human-centered computing Empirical studies in HCIjournalyear: 2024copyright: rightsretainedconference: Companion of the 2024 Computer-Supported Cooperative Work and Social Computing; November 9–13, 2024; San Jose, Costa Ricabooktitle: Companion of the 2024 Computer-Supported Cooperative Work and Social Computing (CSCW Companion ’24), November 9–13, 2024, San Jose, Costa Ricadoi: 10.1145/3678884.3681850isbn: 979-8-4007-1114-5/24/11**footnotetext: These authors contributed equally to this work
@@ 1   Introduction
To address this gap, we propose a LLM-assisted content analysis workflow (Figure 1) that can both save human labor in the analysis and capture the nuances in the content. Recently, LLMs have been tested in various types of qualitative data analysis (Shen et al., 2023), such as deductive thematic analysis (Zhang et al., 2023b); research findings show that LLMs are able to improve coding efficiency (Chew et al., 2023), facilitate collaborative coding (Gao et al., 2023), and increase inter-coder reliability (Xiao et al., 2023). However, most of the research is based on text data. In this paper, we join the exploration of LLMs for data analysis and situate it in the video-data context. Using LLMs, we are specifically interested in customizing codebooks and designing prompts to comprehend and capture more fine-grained information to encompass the semantic and contextual information conveyed in the video content.
@@ 2   Context and Dataset: Depression-
We plan to use LLMs to analyze the content of YouTube videos related to users’ self-disclosure of depression. Online self-disclosure, defined as sharing sensitive personal information in an online public space (Misoch, 2015; Andalibi et al., 2017), has been extensively studied on text-based platforms such as Twitter (Lachmar et al., 2017), Weibo (Wang et al., 2019), and Reddit (Choudhury and De, 2014). The multimodal nature of video may enable deeper levels of disclosure and facilitate a stronger sense of connection between creators and their audience (Liu and Zhang, 2023, 2024). By applying LLMs to extract quantifiable information from the videos automatically, we aim to enable future studies of exploring how people convey information and emotion via self-disclosure videos (Chaudoir and Fisher, 2010). This case study only uses this flow as a preliminary method to analyze video content to explore the capability and limitations of LLMs in content analysis.
@@ 5.1. Future Opportunities and Challenges
LLMs could reduce the human labor required for manual coding in video content analysis. Researchers can create a structured framework (e.g., a codebook) to guide LLMs to automatically identify and extract meaningful insights from unstructured video data such as User Created Content (UGC) in video-sharing platforms and lab/field recordings and eye-tracking data in contexts such as Human-Robot Interactions and Virtual Reality (Hauser et al., 2024; Zhang et al., 2023a). In our case, operationalizing the concept of depth of self-disclosure into various observable behaviors in different settings, such as the presence of other people and the behaviors disclosed, allows us to understand the nuances of how individuals express themselves in video content. However, we find the most critical step in this process is operationalizing concepts into well-defined codes, as it sets the foundation for accurate and comprehensive analysis by LLMs. Further, it is difficult for researchers to curate prompts to harness the power of LLMs (e.g., annotating abstract concepts such as genre in images, as shown in our case study) and effectively evaluate the structured output generated by LLMs. Future work can investigate how to better involve humans in the workflow.